# A comprehensive review of assistive technologies for children with dyslexia

### v1

@@ 7. Research Questions
7.1. RQ1:What are the current interventions that enhance the learning experience for children with dyslexia, and what are their key features?
@@
7.2. RQ2:What are the common accessibility features integrated into these interventions for children with dyslexia?
Taking both technological and non-computing devices into account most of the applications targeted the common challenge of dyslexia. This includes enhancing phonological awareness and letter-sound recognition. Immersive technologies like Virtual reality and augmented reality were focused on gamified activities that involved object’s and shape identification which were useful training for adhering attention. Non-computing devices provided learning platform, engaging with physical objects like books, colors, cards, whiteboard, it also contains structural exercise to bolster active learning. Different word games like syllabic identification, exercise related to identifying consonant and vowels sounds, spelling test, etc. were common in both technological and non-compositional devices. Many mobile applications were developed, with word games, taking account to the strategies of non-computing methodologies, thereby filling the gap it posses. Many technological devices presented audio cues as assistive features to enhance reading fluency and accuracy, similarly OCR and story based app were common in enhancing readability.
@@
7.3. RQ3: What gaps exist in current assistive technology research for children with dyslexia, and what opportunities exist for innovation?
@@
There were limited application which provided user’s personalization, hence personalized application could be the next domain to explore at. Despite these gaps, there are also some opportunities for innovation in this field to consider. First in-cooperating haptics feedback on mobile devices, secondly the use of LLM or NLP technology supporting to comprehend complex sentence in a more nuanced and simplified version. Lastly I further see potential in integration of AI on creating adaptive learning tools. AI could enable real-time analysis of a child’s performance and dynamically adjust tasks to tailor their skill level, ensuring personalized and effective learning experiences.
8. Discussion
In an exploration of the detailed literature review, I was fascinated by the use case of augmented reality in addressing the needs of dyslexic children. AR application are functional on most mobile or tablet devices, and popular being portable and affordable. The application provides 3D animated graphics, which make the learning environment more active, engaging and entertaining. AR are highly advance interm of functionality and popular for it’s mixed realities.
Although existing technologies for children with dyslexia uses smart and advance functionalities they possess some limitations. While, assistive technologies are expected to be both effective and affordable modern technological devices such as virtual reality, and haptic are expensive and complex, which might limit the reach of most children with dyslexia. Virtual device can further exacerbate eye strain with increased screen time. Likewise, growing children experience significant physiological changes, leading to fluctuating needs that can quickly render educational software obsolete or ineffective So, this barrier demands for more personalized applications that can solve the problem that growing children might face.
@@ 8. Discussion
9. Conclusion
The collective findings from the papers on assistive technologies for children with dyslexia highlights the importance of personalized applications to address unique learning challenges faced by children. Moreover, emerging technologies such as virtual reality, augmented reality, and haptic, tangible user interface has demonstrated a significant potential in enhancing and motivating children with learning disabilities. These technologies can provide a more realistic solution by incorporating all (VAKT), a multi-sensory model. Likewise, non-computing strategies, have proven to be efficient in reinforcing memory, for enhancing phonological awareness and can also be used as alternative method of learning where digital technologies are inaccessible.
Acknowledgements.
This Systematic Literature Review was completed as a part of my coursework. I extend my sincere gratitude to Professor Krishna Venkatasubramanian, for his guidance, support, and valuable insights throughout the semester.
Appendix A Appendix
List of abbreviations and acronyms:
S.N Abbreviation Description
1 3D Three Dimensional
2 SLR Systematic literature Review
3 AR Augmentative Reality
4 TUI Tangible User Interface
5 VR Virtual Reality
6 OCR Optical Character Recognition
7 VAKT Visual, Auditory, Kinesthetic, and Tactile
Table 8. List of Abbreviations and Descriptions

### v2

@@ 7 Research Questions
7.1 RQ1:What are the current interventions that enhance the learning experience for children with dyslexia, and what are their key features?
@@
7.2 RQ2:What are the common accessibility features integrated into these interventions for children with dyslexia?
Taking both technological and non-computing devices into account most of the applications targeted the common challenge of dyslexia. This includes enhancing phonological awareness and letter-sound recognition. Immersive technologies like VR and AR were focused on gamified activities that involved object’s and shape identification which were useful training for adhering attention. Non-computing devices provided learning platform, engaging with physical objects like books, colors, cards, whiteboard, it also contains structural exercise to bolster active learning. Different word games like syllabic identification, exercise related to identifying consonant and vowels sounds, spelling test, etc. were common in both technological and non-compositional devices. Many mobile applications were developed, with word games, taking account to the strategies of non-computing methodologies, thereby filling the gap it posses. Many technological devices presented audio cues as assistive features to enhance reading fluency and accuracy, similarly OCR and story based app were common in enhancing readability.
@@
7.3 RQ3: What gaps exist in current assistive technology research for children with dyslexia, and what opportunities exist for innovation?
@@
There are limited applications that offered user personalization, making personalized applications the next domain to explore. Despite these gaps, there are also some opportunities for innovation in this field to consider. First, the integration of haptic feedback into mobile interfaces may enrich user interaction by providing tactile responses, thus improving engagement and accessibility. Second, the incorporation of LLMs and NLP techniques holds potential for improving the comprehension of complex sentences through context-aware simplification and semantic parsing. Lastly, the application of AI in adaptive learning environments presents a compelling direction. AI could enable real-time analysis of a child’s performance and dynamically adjust tasks to tailor their skill level, ensuring personalized and effective learning experiences.
8 Discussion
The systematic literature review highlights the promising role of AR in supporting children with dyslexia. AR applications, which are widely compatible with mobile and tablet devices, are particularly advantageous due to their portability, affordability, and accessibility. These applications often incorporate 3D animated graphics and interactive content, creating immersive, engaging, and multisensory learning environments that cater well to the diverse needs of dyslexic learners
Despite the advancements in assistive technologies, several limitations persist. While tools such as VR and haptic feedback systems demonstrate strong potential for enhancing learning outcomes, their high cost and technical complexity may restrict widespread adoption, especially in low-resource settings. Furthermore, VR systems can contribute to increased screen time, potentially causing eye strain or fatigue, particularly among younger users. As children undergo rapid developmental changes, their learning needs may evolve, which can render static or non-adaptive educational tools less effective over time. These factors underscore the necessity for personalized, adaptable interventions that evolve alongside the child’s cognitive and physiological growth
@@ 8 Discussion
Another significant challenge lies in the design of assistive applications themselves. Tools created specifically for learners with disabilities often feature distinct interface designs—such as enlarged fonts or high-contrast colors—which, while functional, may inadvertently draw attention to the user’s differences. This can contribute to social stigma, ultimately discouraging children from consistently using such tools. Therefore, ensuring that assistive technologies are discreet, inclusive, and socially acceptable is critical for long-term adoption and efficacy. Our work directs researchers or developers on creating more personalized, cross platform technologies in-cooperating multi-sensory modalities for recreating enhancing technologies.
9 Conclusion
The collective findings from the papers on assistive technologies for children with dyslexia highlights the importance of personalized assistive technologies in addressing the unique learning challenges faced by children with dyslexia. Emerging tools such as VR, AR, haptics, and TUI have shown strong potential to enhance motivation and learning outcomes through multisensory (VAKT) engagement. These technologies offer more interactive and tailored experiences. Additionally, non-computing strategies remain effective in improving phonological awareness and memory, serving as valuable alternatives in contexts with limited access to digital resources.
[…]
Acknowledgements
We would like to express sincere gratitude to Professor Krishna Venkatasubramanian of the University of Rhode Island for his valuable guidance guidance, constructive feedback, and continuous support throughout the development of this work.
10 Disclosure of Interest
The authors declare that there are no known competing interests.
11 Disclosure of funding statement
Authors declare that they received no financial aid or funding for this work.
12 Disclosure of Data
This study did not generate any new datasets. All data analyzed in this review were obtained from previously published studies, which are properly cited within the manuscript.